A sample containing 33.42g of metal pellets is poured into a graduated cylinder initially containing 12.7ml of water , causing the water level in the cylinder to rise to 21.6 ml Calculate the density of the metal .
1
Expert's answer
2016-11-04T11:45:10-0400
Volume of the metal: 21.6 – 12.7 = 8.9 ml (cm3)
Empirical formula of density:
ρ=m/V
ρ(metal)=33.42/8.9 = 3.8 g/cm3